OpenStudy (anonymous):

the derivative of 6t^-3 would be 0 because i got 0 but i plug it in my online hw and says it's incorrect

OpenStudy (anonymous):

is this for the hw due tonight? hah

OpenStudy (anonymous):

smh yea man it is and i have like 3 problems left to do?

OpenStudy (anonymous):

I have like 7, and 40 mins to finish it. What problem are you on?

OpenStudy (anonymous):

i have problem 6, 14 , 17 , and 19 to do.

OpenStudy (anonymous):

and the derivative is -18/(t^4) if you wrote that function down correctly

OpenStudy (anonymous):

it is how does the power keep going down for the negative ones?

OpenStudy (anonymous):

I'd explain if I wasn't pressed for time. Just plug it in and get it over with for now.

OpenStudy (anonymous):

hahah aiight man i just did, it worked thanks man. ima let you get back and finish before time is up

OpenStudy (anonymous):

for problem 2 and those alike, how do you find f '(3) or whatever it's asking?

OpenStudy (anonymous):

plug in the derivative equation you got into the calculator and look for the f ' of 3 and that should be your answer

OpenStudy (anonymous):

still lost. my f(x) is 5+3/x+4/x^2. the f'(x) is -((3x+8)/(x^3)). it's asking for f '(3)

OpenStudy (anonymous):

plug in the f'(x) is -((3x+8)/(x^3)) into the calculator and go to table and look what f'(3) is, that should be your answer.

OpenStudy (anonymous):

i'm retarded, I had my symbols messed up :/ worked, thanks

OpenStudy (anonymous):

haha its all good man, it happened to me with switching pos. and neg. no problem

OpenStudy (anonymous):

have you worked out problem 14 yet, its confusing?

OpenStudy (anonymous):

yeah, what's your f(x)?

OpenStudy (anonymous):

and what values is it asking for?

OpenStudy (anonymous):

and how did you do problem 8?

OpenStudy (anonymous):

its f(x)= 5/5sqrt(X)

OpenStudy (anonymous):

5*sqrt(x) or x^(1/5)?

OpenStudy (anonymous):

the 2nd one, i didnt know how to write it out

OpenStudy (anonymous):

is it asking for f '(1) and f '(5)?

OpenStudy (anonymous):

yea its asking for both of them

OpenStudy (anonymous):

-1 and -.145 respectively

OpenStudy (anonymous):

thanks. man

OpenStudy (anonymous):

for problem 8 you have to find the derivative and then put the derivative into the calulator and find x where y is 0.

OpenStudy (anonymous):

there should be two solutions

OpenStudy (anonymous):

my derivative is 30x^2+90x-12. can't find them

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

no wait. 30x^2+90x-120

OpenStudy (anonymous):

nvm, got it.

OpenStudy (anonymous):

the answer should be -4 and 1.

OpenStudy (anonymous):

are you done? I need just need to get 19 and the last part for 16.

OpenStudy (anonymous):

naw i need to do 17 and 19 too.

OpenStudy (anonymous):

v=s'=4t^3-3t^2 a=12t^2-6t=0 2t^2-t=0 t(2t-1)=0 t=0 or t=1/2 was what I was given for 17. your equation might be different though

OpenStudy (anonymous):

for the last part of 16 the answer for M: would be the same as the derivative of f'(4) and to get B you have to use the formula y-y1=m(x-x1) which will give you the whole equation in y=mx+b form

OpenStudy (anonymous):

i mean whatever derivative you got

OpenStudy (anonymous):

i'll do your 17 if you do my 16.

OpenStudy (anonymous):

my f(x) = 4x+2/x. f '(5) = (4x^2-2)/x^2 => 3.92, which was also m. I just need b

OpenStudy (anonymous):

A particle moves along a straight line with equation of motion s=t^3-3t^2.

OpenStudy (anonymous):

try 1.

OpenStudy (anonymous):

your first derivative is 3t^2-6t, and your second is 6t-6 => 6t-6=0 6t=6 t=6/6

OpenStudy (anonymous):

m: 3.92 b: .8

OpenStudy (anonymous):

.8 worked, cheers.

OpenStudy (anonymous):

y-y1=m(x-x1) y-20.4=3.92(x-5) y-20.4=3.92x-19.6 y=3.92x+.8

OpenStudy (anonymous):

1 worked tooooo man finally done thanks a lot man!!

OpenStudy (anonymous):

did you get 19? I have 7 mins to finish that pelletttt

OpenStudy (anonymous):

mine's find d^2*y/dx^2 for y=-5x^4+4x

OpenStudy (anonymous):

naw man thats the only one im having trouble on...i think you should ask someone!?!

OpenStudy (anonymous):

i did, like 10 times but no one has answered so far

OpenStudy (anonymous):

what should we do..i think the professor will open it 2morrow if we ask him cause he has done tht for all the other ones we can ask him to work tht problem

OpenStudy (anonymous):

it has the step by step?

OpenStudy (anonymous):

it's not the same as our though

OpenStudy (anonymous):

yea i think we can work with it, doesn't seem to have a lot of steps

OpenStudy (anonymous):

2 mins, good luck

OpenStudy (anonymous):

haha ima put it into wolf ram alpha nd get the answer

OpenStudy (anonymous):

i got it what's your eq?

OpenStudy (anonymous):

mine is y=-8x^4+4x

OpenStudy (anonymous):

-96x^2

OpenStudy (anonymous):

got it right and im done!!!!

OpenStudy (anonymous):

lol same, what a pain... who are you, by the way?

OpenStudy (anonymous):

gabriel the dude with the snapback, umm we stayed after school with the other guy to finish hw

OpenStudy (anonymous):

oh right, I told you about this site. lol if we're relying on this for our hw then we're gonna mess up on that test big time

OpenStudy (anonymous):

hahaha..yea i try only to ask 3 questions or fewer....the others i had to look at youtube videos and they really helped

OpenStudy (anonymous):

we have review all class 2morrow so we gonna be set

OpenStudy (anonymous):

yeah, but with summer who has time to be actually studying and pellet? did you get the book?

OpenStudy (anonymous):

i know i be tired after tht 3 hour lecture man its coming on wednesday..bout to be in it for help and understanding

OpenStudy (anonymous):

hope it helps you man. Aight, I'm gonna go shower and pass out. see you in class

OpenStudy (anonymous):

aiight man see man